# SOC-DGL: Social Interaction Behavior Inspired Dual Graph Learning Framework for Drug-Target Interaction Identification

### v1

@@ 6.5 Validation experiment on imbalanced loss function in imbalanced datasets
The standard binary cross-entropy loss function was a commonly used loss function suitable for balanced datasets, but it might have caused the model to favor the majority class in imbalanced datasets. The focal loss [51] aimed to address class imbalance by reducing the weight of easily classified samples while increasing the weight of hard-to-classify samples. The weighted binary cross-entropy loss function balanced the disparity between positive and negative samples by assigning higher weights to the minority class. The reduced negative weighted binary cross-entropy loss function reduced the weight of negative samples while keeping the weight of positive samples unchanged, mitigating the impact of imbalanced data on model performance. The formulas for the four loss functions were as follows:
@@
$\mathcal{L}_{F}=-\frac{|\mathcal{Y}^{-}|}{|\mathcal{Y}^{+}|}\left(1-h_{ij}^{*}\right)^{\gamma}\log(h_{ij}^{*})$ (28)
$\mathcal{L}_{W}=-\frac{1}{n_{d}\cdot n_{t}}(\frac{\mid y^{-}\mid}{\mid y^{+}\mid}\!\!\sum_{(i,j)\in y^{+}}\!\!\!\!\log(h_{ij}^{*})+\!\!\!\!\sum_{(i,j)\in y^{-}}\!\!\!\!\log(1-h_{ij}^{*}))$ (29)
$\mathcal{L}_{R}=-\frac{1}{n_{d}\cdot n_{t}}(\varpi\frac{\mid y^{-}\mid}{\mid y^{+}\mid}\!\!\sum_{(i,j)\in y^{+}}\!\!\!\!\log(h_{ij}^{*})+\!\!\!\!\sum_{(i,j)\in y^{-}}\!\!\!\!\log(1-h_{ij}^{*}))$ (30)
where $\mathcal{L}_{S}$, $\mathcal{L}_{F}$, $\mathcal{L}_{W}$, and $\mathcal{L}_{R}$ were Standard Binary Cross-Entropy Loss (SLF), Focal Loss with Dynamic Scaling (FLF), Weighted Binary Cross-Entropy Loss (WLF), and Weighted Binary Cross-Entropy Loss with Reduced Negative Samples (RLF), respectively. $n_{d}$ was the number of drugs in the dataset, $n_{t}$ was the number of targets in the dataset, $|y^{+}|$ represented the number of positive samples, and $|y^{-}|$ represented the number of negative samples. $h_{ij}^{*}$ represented the model’s predicted probability of interaction between drug $i$ and target $j$. $\gamma$ was an adjustment parameter used to reduce the weight of easily classified samples and increase the weight of hard-to-classify samples. $\varpi$ was also an adjustment parameter, used to reduce the weight of negative samples while keeping the weight of positive samples unchanged.

### v2

@@ 6.5 Validation experiment on imbalanced loss function in imbalanced datasets
The formulas for the four loss functions (SLF, FLF, WLF, and RLF) are as follows:
@@
$\mathcal{L}_{F}=-\frac{|y^{-}|}{|y^{+}|}\left(1-h_{ij}^{*}\right)^{\gamma}\log(h_{ij}^{*})$ (28)
$\mathcal{L}_{W}=-\frac{1}{n_{d}\cdot n_{t}}(\frac{|y^{-}|}{|y^{+}|}\!\!\sum_{(i,j)\in y^{+}}\!\!\!\!\log(h_{ij}^{*})+\!\!\!\!\sum_{(i,j)\in y^{-}}\!\!\!\!\log(1-h_{ij}^{*}))$ (29)
$\mathcal{L}_{R}=-\frac{1}{n_{d}\cdot n_{t}}(\varpi\frac{|y^{-}|}{|y^{+}|}\!\!\sum_{(i,j)\in y^{+}}\!\!\!\!\log(h_{ij}^{*})+\!\!\!\!\sum_{(i,j)\in y^{-}}\!\!\!\!\log(1-h_{ij}^{*}))$ (30)
where $\mathcal{L}_{S}$, $\mathcal{L}_{F}$, $\mathcal{L}_{W}$, and $\mathcal{L}_{R}$ represent the SLF, FLF, WLF, and RLF, respectively. $n_{d}$ denotes the number of drugs in the dataset, $n_{t}$ is the number of targets, $|y^{+}|$ is the count of positive samples, and $y^{-}|$ is the count of negative samples. $h_{ij}^{*}$ represents the predicted interaction probability between drug $i$ and target $j$. The adjustment parameter $\gamma$ modulates the weight of easily classified samples, and $\varpi$ adjusts the weight of negative samples without affecting positive sample weights.